# Unfitted finite element modelling of surface-bulk viscous flows in animal cells

### v1

@@ Acknowledgments
We thank all members of the Turlier team for the fruitful discussions and Gridap developers for the many contributions that made this work possible. E.N. received funding from the European Union’s Horizon EU research and innovation programme under grant agreement no. 101105565. H.T. received funding from the European Union’s Horizon 2020 research and innovation programme under the European Research Council grant agreement no. 949267 and has been supported by the Bettencourt-Schueller Foundation, by the CNRS and the Collège de France.
@@ Appendix A Axisymmetric formulation in 2D
Given the differential operators in cylindrical coordinates $(r,\theta,z)$ [Bernardi1999], in particular, for a scalar field $v$ and a vector field $\boldsymbol{u}$,
~$\tilde{\nabla}v=\nabla v,\enskip\tilde{\nabla}\cdot u=\frac{1}{r}\nabla\cdot(r\boldsymbol{u}),\enskip\text{and}\enskip\tilde{\boldsymbol{\varepsilon}}(\boldsymbol{u})=\begin{bmatrix}\frac{\partial u^{r}}{\partial r}&0&\frac{1}{2}\left(\frac{\partial u^{r}}{\partial z}+\frac{\partial u^{z}}{\partial r}\right)\\ 0&\frac{u^{r}}{r}&0\\ \frac{1}{2}\left(\frac{\partial u^{r}}{\partial z}+\frac{\partial u^{z}}{\partial r}\right)&0&\frac{\partial u^{z}}{\partial z}\\ \end{bmatrix},$

### v2

@@ Acknowledgments
We thank all members of the Turlier team for the fruitful discussions and Gridap developers for the many contributions that made this work possible. E.N. received funding from the European Union’s Horizon EU research and innovation programme under grant agreement no. 101105565. H.T. received funding from the European Union’s Horizon 2020 research and innovation programme under the European Research Council grant agreement no. 949267, from the Human Frontier Science Program research grant no. RGP023/2025 and has been supported by the CNRS and the Collège de France.
@@ Appendix A Axisymmetric formulation in 2D
Given the differential operators in cylindrical coordinates $(r,\theta,z)$ [21], in particular, for a scalar field $v$ and a vector field $\boldsymbol{u}$,
$\tilde{\nabla}v=\nabla v,\enskip\tilde{\nabla}\cdot u=\frac{1}{r}\nabla\cdot(r\boldsymbol{u}),\enskip\text{and}\enskip\tilde{\boldsymbol{\varepsilon}}(\boldsymbol{u})=\begin{bmatrix}\frac{\partial u^{r}}{\partial r}&0&\frac{1}{2}\left(\frac{\partial u^{r}}{\partial z}+\frac{\partial u^{z}}{\partial r}\right)\\ 0&\frac{u^{r}}{r}&0\\ \frac{1}{2}\left(\frac{\partial u^{r}}{\partial z}+\frac{\partial u^{z}}{\partial r}\right)&0&\frac{\partial u^{z}}{\partial z}\\ \end{bmatrix}$,